A certain gasoline weighs 46.5 lb/ft3 . What are its mass density, specific volume and specific gravity?


1
Expert's answer
2022-02-14T12:58:21-0500

"1 \:\rm lb/ft^3 = 16.018\:kg/m^3"

"\rho=46.5*\rm 16.018\:kg/m^3=745\:kg/m^3"

"1/\rho=1/745\:\rm kg/m^3=0.0013\: m^3/kg"

"\rho/\rho_w=745\:\rm kg/m^3/1000\: kg/m^3=0.745"